OpenStudy (anonymous):

Suppose y varies directly with x. If x = 7 when y = 23, what is the value of y when x = 21? Select one: a. 59 b. 69 c. 89 d. 79

OpenStudy (anonymous):

7/23 = 21/x cross multiply (7)(x) = (21)(23) 7x = 483 x = 483/7 x = 69 answer is 69

OpenStudy (anonymous):

Thanks :) !